OpenStudy (anonymous):

find the integral of (sin(1/x)) dx

OpenStudy (lgbasallote):

\[\huge \int \sin \left(\frac 1x \right)dx\] this is the question yes?

OpenStudy (anonymous):

yes..

OpenStudy (lgbasallote):

this doesnt seem simple

OpenStudy (mimi_x3):

integral that does not have a closed form? http://www.wolframalpha.com/input/?i=integrate++%28sin%281%2Fx%29%29+dx

OpenStudy (anonymous):

hmmm.. im not understand form the answer what's means of Ci?

OpenStudy (mimi_x3):

is it by chance a definite integral?

OpenStudy (anonymous):

ok, thanks yea

OpenStudy (mimi_x3):

if its a definite integral i think you can do it numerically

OpenStudy (anonymous):

no closed form for it but u can represent it by series